OpenStudy (anonymous):

Write an indirect proof proving the two non-right angles of a right triangle measure less than 90°.

OpenStudy (welshfella):

but they don't..

OpenStudy (anonymous):

I don't ask questions, I just need the answers. Can you pull anything out of this for an answer?

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{\normalsize \text{let's consider that in right angled triangle ABC}\hspace{.33em}\\~\\ \normalsize \text{A and B are non right angles and C is right angle }\hspace{.33em}\\~\\ \normalsize \text{let us assume A+B>90 }\hspace{.33em}\\~\\ and~~ C=90 \hspace{.33em}\\~\\ A+B+C=180 \hspace{.33em}\\~\\ A+B+90=180 \hspace{.33em}\\~\\ A+B=90 \hspace{.33em}\\~\\ \normalsize \text{but }\hspace{.33em}\\~\\ A+B>90 \hspace{.33em}\\~\\ \normalsize \text{hence this contradicts our idea that A+b>90 }\hspace{.33em}\\~\\ \normalsize \text{hence } A+B= 90\hspace{.33em}\\~\\ \normalsize \text{hence } A< 90\hspace{.33em}\\~\\ \normalsize \text{hence } B<90\hspace{.33em}\\~\\ }\end{align}\)